OpenStudy (anonymous):

Find the y-coordinate of the vertex of the parabola whose equation is y = x^2 - x + 2. A.) 1 3/4 B.) 2 C.) 2 1/4

OpenStudy (anonymous):

I'd say it's 2, because if you're following the equation y=a(x-h)+k then the k would be the y-coordinate of the vertex....and the x would be 0 and so the h(x) would be the x-coordinate in the vertex

OpenStudy (shamil98):

Set x = 0 ..

OpenStudy (shamil98):

y = 0^2 - 0 + 2 = ?